My room is 242 inches long and 134 inches wide how many square feet is it?
Answer by jim_thompson5910(35256)

First convert from inches to feet
242 inches = (242 inches)*(1 ft/12 inches) = (242/12) ft = (121/6) ft
134 inches = (134 inches)*(1 ft/12 inches) = (134/12) ft = (67/6) ft

Summary of conversions
Length = 242 inches = (121/6) ft
Width = 134 inches = (67/6) ft

-----------------------------------------------------------------

Area of rectangle = (Length) * (Width)
Area of rectangle = (121/6 ft) * (67/6 ft)
Area of rectangle = ((121/6)*(67/6))*(ft*ft)
Area of rectangle = ((121*67)/(6*6))*(ft^2)
Area of rectangle = (8107/36)*(ft^2)
Area of rectangle = (225.194444)*(ft^2)
Area of rectangle = 225.194444 ft^2

The notation "ft^2" means "square feet"

The area of the floor is approximately 225.194444 square feet.